OpenStudy (anonymous):

what is ((x-4)/6) ÷ ((x+1)/7)

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

I will fan and medal

OpenStudy (anonymous):

do u want the answer or help

OpenStudy (anonymous):

both if you don't mind

OpenStudy (anonymous):

ok so the answer is 6

OpenStudy (anonymous):

and heres how to do it its long so get ready

OpenStudy (anonymous):

Simplifying -7 + x + 4 + 4x = x + -3 + 6 + 3x Reorder the terms: -7 + 4 + x + 4x = x + -3 + 6 + 3x Combine like terms: -7 + 4 = -3 -3 + x + 4x = x + -3 + 6 + 3x Combine like terms: x + 4x = 5x -3 + 5x = x + -3 + 6 + 3x Reorder the terms: -3 + 5x = -3 + 6 + x + 3x Combine like terms: -3 + 6 = 3 -3 + 5x = 3 + x + 3x Combine like terms: x + 3x = 4x -3 + 5x = 3 + 4x Solving -3 + 5x = 3 + 4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4x' to each side of the equation. -3 + 5x + -4x = 3 + 4x + -4x Combine like terms: 5x + -4x = 1x -3 + 1x = 3 + 4x + -4x Combine like terms: 4x + -4x = 0 -3 + 1x = 3 + 0 -3 + 1x = 3 Add '3' to each side of the equation. -3 + 3 + 1x = 3 + 3 Combine like terms: -3 + 3 = 0 0 + 1x = 3 + 3 1x = 3 + 3 Combine like terms: 3 + 3 = 6 1x = 6 Divide each side by '1'. x = 6 Simplifying x = 6

OpenStudy (anonymous):

ok so my choices were -1 4 -7 none of the above

OpenStudy (anonymous):

so it is d

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

glad i could help

OpenStudy (anonymous):

it was A

OpenStudy (anonymous):

but thanks anyway